Dr.Thomas McKay lived 3/13 of his life as a child.He spent 4/39 of his life preparing for his outstanding medical career.For 1/2 of his life, he was a very successful surgeon in a famous children's hospital.Since his retirement,he has lived 13 years on a comfortable pension.How old is he now?
Answer by rfer(16322)

3/3+4/39+1/2=18/78+8/78+39/78=65/78
----------------------------------------
13/(13/78)=
13*78/13=1014/13=78 yrs